# A survey on the resolvent convergence

### v1

@@ 2 Resolvent and spectrum of an operator
    If the inverse of $A-\lambda$ exists, then it is of course linear, and by the closed graph Theorem (ReedSimon1980, , Theorem III.12) it is also bounded. Hence, the boundedness requirement for the inverse in the definition is redundant. However, the resolvent set uses to appear in the literature defined like this; see (ReedSimon1980, , Definition in pg. 253) or (Teschl2014, , Section 2.4).
@@ 3 Resolvent convergence
$\mathrm{Dom}(P)$ $:=$ $\{\varphi\in L^{2}(\mathbb{R}):\int_{\mathbb{R}}x^{2}|\varphi(x)|^{2}\,dx<+\infty\},$
@@ 4 Convergence of spectra
We revisit Example 1, where we have shown that, in $H=L^{2}(\mathbb{R})$, the scaled position operator $A_{n}:=\frac{1}{n}P$ converges to the zero operator in the strong resolvent sense as $n\to+\infty$ (and not in the norm resolvent sense). The limiting operator, $A=0$, of course has trivial spectrum $\sigma(A)=\{0\}$. Instead, we show that $\sigma(P)=\mathbb{R}$, and hence $\sigma(A_{n})=\mathbb{R}$ for every $n\in\mathbb{N}$.
@@ 5 Relation with strong graph limit
Indeed, if an analogous statement to Theorem 5.1 was true for the weak graph limit, then by Theorem 5.2 we would have that, for uniformly bounded self-adjoint operators, weak resolvent convergence would be equivalent to convergence in the weak topology, which contradicts Corollary 2. This proves the following statement.
@@ 6 Relation with $G$- and $\Gamma$-convergence
With an abuse of notation, we alternatively say that $Q_{n}$ strongly $\Gamma$-converges to $Q_{\infty}$, and we write $Q_{n}\overset{\mathrm{S\Gamma}}{\underset{n\to+\infty}{\longrightarrow}}Q_{\infty}$. Assuming, in addition, that there exists a constant $c>0$ such that $Q_{n}(v)\geq c\|v\|_{H}^{2}$ for every $v\in H$ and every $n\in\mathbb{N}\cup\{\infty\}$, if in Definition 9 we replace the strong convergence of $u_{n}$ by weak convergence, then we get a characterization of the $\Gamma$-convergence in the weak topology of $Q_{n}$ to $Q_{\infty}$; see (DalMaso1993, , Chapter 8). With an abuse of notation, we alternatively say that $Q_{n}$ weakly $\Gamma$-converges to $Q_{\infty}$, and we write $Q_{n}\overset{\mathrm{W\Gamma}}{\underset{n\to+\infty}{\longrightarrow}}Q_{\infty}$.
Similarly, the notion of $G$-convergence of coercive self-adjoint operators is the same as in (DalMaso1993, , Definition 13.3).
@@
The reader can find in Section 7 a diagram summarizing the results involving the resolvent convergence that we stated in this survey; see Figure 1.
@@ Acknowledgements.
The author thanks Maicol Caponi, Alessandro Carbotti, Giuseppe Cosma Brusca, and Alberto Maione for enlightening discussions regarding the relations between resolvent and $G$- and $\Gamma$-convergence.

### v2

@@ 2 Resolvent and spectrum of an operator
    If the inverse of $A-\lambda$ exists, then it is linear, and by the closed graph Theorem (ReedSimon1980, , Theorem III.12) it is also bounded. Hence, the boundedness requirement for the inverse in the definition is redundant. However, the resolvent set uses to appear in the literature defined like this; see (ReedSimon1980, , Definition in pg. 253) or (Teschl2014, , Section 2.4).
@@ 3 Resolvent convergence
$\mathrm{Dom}(P)$ $:=$ $\left\{\varphi\in L^{2}(\mathbb{R}):\int_{\mathbb{R}}x^{2}|\varphi(x)|^{2}\,dx<+\infty\right\},$
@@ 4 Convergence of spectra
We revisit Example 1, where we have shown that, in $H=L^{2}(\mathbb{R})$, the scaled position operator $A_{n}:=\frac{1}{n}P$ converges to the zero operator in the strong resolvent sense as $n\to+\infty$ (and not in the norm resolvent sense). The limiting operator, $A=0$, has trivial spectrum $\sigma(A)=\{0\}$. Instead, we show that $\sigma(P)=\mathbb{R}$, and hence $\sigma(A_{n})=\mathbb{R}$ for every $n\in\mathbb{N}$.
@@ 5 Relation with strong graph limit
Indeed, if an analogous statement to Theorem 5.1 were true for the weak graph limit, then by Theorem 5.2 we would have that, for uniformly bounded self-adjoint operators, weak resolvent convergence would be equivalent to convergence in the weak topology, which contradicts Corollary 2. This proves the following statement.
@@ 6 Relation with $G$- and $\Gamma$-convergence
Two remarks are in order about Definition 9.
  1. 1.
    With an abuse of notation, we alternatively say that $Q_{n}$ strongly $\Gamma$-converges to $Q_{\infty}$, and we write $Q_{n}\overset{\mathrm{S\Gamma}}{\underset{n\to+\infty}{\longrightarrow}}Q_{\infty}$.
  2. 2.
    There exists a notion of “weak” $\Gamma$-convergence, the definition of which is not analogous to Definition 9; see (Braides2002, , equation (1.13) and the paragraph below it). However, assuming, in addition, that there exists a constant $c>0$ such that $Q_{n}(v)\geq c\|v\|_{H}^{2}$ for every $v\in H$ and every $n\in\mathbb{N}\cup\{\infty\}$, if in Definition 9 we replace the strong convergence of $u_{n}$ by weak convergence, then we get a characterization of the so-called $\Gamma$-convergence in the weak topology of $Q_{n}$ to $Q_{\infty}$; see (DalMaso1993, , Proposition 8.10). With an abuse of notation, we alternatively say that $Q_{n}$ weakly $\Gamma$-converges to $Q_{\infty}$, and we write $Q_{n}\overset{\mathrm{W\Gamma}}{\underset{n\to+\infty}{\longrightarrow}}Q_{\infty}$.
Similarly, the notion of $G$-convergence of equi-coercive self-adjoint operators is the same as in (DalMaso1993, , Definition 13.3).
@@
It is worth noting that, by Theorem 6.1, the strong resolvent convergence of self-adjoint operators is not equivalent to the strong $\Gamma$-convergence of their quadratic forms. The main reason for this is the fact that, in general, the weak and strong $\Gamma$-limits of a same sequence may differ; see (DalMaso1993, , Example 6.6). The following result provides a scenario in which the strong resolvent convergence is equivalent to all the notions of convergence that have been presented in this section.
Theorem 6.2
Assume that $(h,\langle\cdot,\cdot\rangle_{h},\|\cdot\|_{h})$ is a compactly embedded Hilbert subspace of $H$. Let $A_{n}$ and $A_{\infty}$ be self-adjoint operators in $H$ such that
  1. 1.
    $\mathrm{Dom}(A_{n})\subseteq h$ for every $n\in\mathbb{N}\cup\{\infty\}$, and
  2. 2.
    their associated quadratic forms are equi-coercive in $h$, namely, they satisfy
[…]
Then, the following statements are equivalent.
  1. 1.
    $A_{n}$ converges in the strong resolvent sense to $A_{\infty}$.
  2. 2.
    $A_{n}$ strongly $G$-converges to $A_{\infty}$.
  3. 3.
    $Q_{n}$ strongly $\Gamma$-converges to $Q_{\infty}$.
  4. 4.
    $Q_{n}$ weakly $\Gamma$-converges to $Q_{\infty}$.
  5. 5.
    $A_{n}$ weakly $G$-converges to $A_{\infty}$.
A proof of this theorem can be found in (DalMaso1993, , Theorem 13.12), where the Hilbert spaces are assumed to be real. Nevertheless, the real structure is only used in the proof of (DalMaso1993, , Theorem 13.12) to invoke Theorem 6.1, which also holds in the complex Hilbert space setting.
The scenario described by the hypothesis of Theorem 6.2 is usually found in the context of uniform elliptic PDEs in divergence form. The reader can find one such concrete example in Section 8, where we study the resolvent convergence in $H=L^{2}(\Omega)$ of the Robin Laplacian to the Dirichlet Laplacian as the Robin boundary condition degenerates to the Dirichlet boundary condition; see the notation at the beginning of Section 8. Indeed, the quadratic forms associated to the weak formulations of both the Robin and the Dirichlet Laplacians are equi-coercive in the compactly embedded Hilbert subspace $h=H^{1}(\Omega)$, on which the operators are defined.
The reader can also find in Section 7 a diagram summarizing the results involving the resolvent convergence that we stated in this survey; see Figure 1.
@@ Acknowledgements.
The author thanks Maicol Caponi, Alessandro Carbotti, Giuseppe Cosma Brusca, and Alberto Maione for enlightening discussions regarding the relations between resolvent and $G$- and $\Gamma$-convergence. The author also thanks the suggestions of the anonymous referee, which have enriched Section 6.